OpenStudy (anonymous):

Simplify {(5)^3-(2+3)/(10)

OpenStudy (mathstudent55):

Is this it? Or is the 10 only under 2 + 3? \( \dfrac{ (5)^3-(2+3)}{10} \)

OpenStudy (mathstudent55):

You have an open bracket, but you didn't close it, so it's hard to understand what you meant.

OpenStudy (anonymous):

only under 2+3

OpenStudy (mathstudent55):

Ok. I'll try again. This? \(5^3- \dfrac{2+3}{10}\)

OpenStudy (mathstudent55):

If so, start by cubing the 5. What is 5^3 = ?

OpenStudy (anonymous):

\[5^{3}- (2+3)/10 \] this is what i see

OpenStudy (anonymous):

75

OpenStudy (anonymous):

\[5^3-\frac{2+3}{10}=125-\frac{1}{2}=\frac{249}{2} \]

OpenStudy (mathstudent55):

\(5^5 = 5 \times 5 \times 5 = 125\) \(5^3- \dfrac{2+3}{10}\) \(= 125 - \dfrac{2+3}{10}\) Now we work on the numerator of the fraction: 2 + 3 \(= 125 - \dfrac{5}{10}\) \(=125 - 0.5\) Finally, we subtract: \(= 124.5\)

OpenStudy (anonymous):

the answers are A.12 B.123 and C.7/10 i have the answer i just wanted to see you my teacher got it the answer is A @mathstudent55

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

->{ (5)^3-(2+3)/(10) The closing brace is missing. Where is it to be placed?

OpenStudy (mathstudent55):

@AlanaLeigh14 Notice that the problem we solved does not have any of the answers you listed as choices. Maybe the problem is like this: \(5^3- \dfrac{10}{2 + 3} = 125 - \dfrac{10}{5} = 125 - 2 = 123\)